OpenStudy (anonymous):

Is the square root of 48 rational or irrational?

OpenStudy (anonymous):

Irrational

Directrix (directrix):

Irrational

OpenStudy (agreene):

\[4\sqrt{3}\] irrational.

OpenStudy (anonymous):

oh my god theres alot of people looking at this question!!, and thanks!

OpenStudy (anonymous):

\[48 = 2^{4}\times3\] \[\sqrt{48} = 2^{2}\times \sqrt{3}\] \[\sqrt{48} = 4\sqrt{3}= 4\times1.73205080756.... = irrational\]

OpenStudy (anonymous):

ok now is 0.0125 irrational or rational?

OpenStudy (hoblos):

this one is rational

OpenStudy (anonymous):

k

OpenStudy (anonymous):

The formula d = , where h equals the height, in feet, of the viewer's eyes, estimates the distance d, in miles, to the horizon from the viewer. Find the distance to the horizon for a person who eyes are 6 ft above the ground. 2 mi 81 mi 3 mi 9 mi

OpenStudy (anonymous):

i really need help with this, im not good at it at all

OpenStudy (anonymous):

the formula d = ?

OpenStudy (anonymous):

oh wait

OpenStudy (anonymous):

hold on a sec

OpenStudy (anonymous):

D=\[\sqrt{1.5 h}\]

OpenStudy (anonymous):

3 :P

OpenStudy (anonymous):

thanks!!!!! :) you guys helped me so much! :D thanks

OpenStudy (anonymous):

yw ;)